# Evaluating Time-Dependent Methods and Seasonal Effects in Code Technical Debt Prediction

### v1

@@ 3.5.4 Setting the stage for Machine Learning models
In RQ2 and RQ4, we compare the prediction performance of TSA approaches with ML algorithms. Therefore, we also investigate the ability of ML models to predict Code TD by applying a collection of linear and non-linear ML algorithms with the collected data and compare the results with the time-dependent approaches mentioned before.
@@
The ML methods used for the comparison are selected to provide a wide predictive perspective, as each of them emphasizes different aspects of the data. Likewise, the chosen ML algorithms have been extensively used in the recent literature for their ability to predict software quality features, like TD [12, 9, 10, 44, 45].
@@
Multiple Linear Regression (MLR) is a commonly used and powerful statistical model for analyzing the relationship between a continuous dependent variable and a set of independent explanatory variables. Through this relationship, the model can provide insights on how changes in the independent variables affect the value of the response variable while maintaining the linearity assumption [46].
@@
Stochastic Gradient Descent (SGD) algorithm relies on the iterative decrease of prediction loss by changing the model parameters. Precisely, for each iteration, the SGD algorithm randomly selects a training set to bring randomness into the optimization problem, and the algorithm continues optimizing the result until it converges [47].
@@
Lasso regression (L1) is a linear regression technique specialized in high dimensional explanatory variable sets. This technique performs an optimization problem where model-independent variables get penalized if they are not important for the model. Thus, through model simplification, L1 achieves linear regression with the most relevant explanatory variables [48].
@@
Ridge regression (L2) is another linear regression technique that, through a regularization term, avoids overfitting the model. The L2 technique brings a penalization for the given model parameters that denote extreme values and thus stabilizes the model which results in advantages for applications where the number of model-independent variables is high [49].
The selected ML models assuming non-linearity (NL) are:
@@
Support Vector Machine (SVM) is a well-established ML algorithm used for classification and regression tasks. The regression model aims to find the most optimal hyperplane that considers the characteristics of the existing explanatory variables for explaining the values of the output response variable. For this, the model performs a minimization process of the deviation between the real value and the predicted value [50]. We considered using the so-called Gaussian kernel or Radial Basis Function as the function to build the model to get the perspective of a non-linear approach.
@@
Extreme Gradient Boost (XGB) is a ML algorithm part of the gradient boosting method family. Based on the ensemble learning methodology, the XGB algorithm employs multiple decision trees in a greedy format, where each of the subsequent trees corrects or refines the wrong results from the previous tree. Through this technique, the algorithm defines weights for the included independent variables that require a higher emphasis [51].
@@
Random Forest regression (RF) also belongs to the ensemble learning family, and it is built upon the basis of the ensemble of decision trees. Based on random sampling, the RF algorithm performs multiple decision trees and the output is the aggregated value from the results obtained in the performed trees [52].
3.5.5 Performance metrics
@@
We used four performance metrics to compare and evaluate the prediction approaches and models.
@@
The first performance metric is the Mean Absolute Percentage Error (MAPE). The MAPE is a statistical metric commonly used for measuring prediction accuracy. It quantifies the magnitude of errors between the predicted values and the actual observed values, calculating the mean of the absolute value of the prediction error. Its formula is:
@@
where is the number of observations, is the i-th actual observed value and is the i-th predicted value. As shown by the formula, MAPE is represented as a percentage metric, where a smaller percentage of error means better predictive performance and the opposite with a higher percentage.
@@
MAPE has limitations when the actual observations are small or close to zero, thus introducing bias into the model training. Therefore, we selected alternative accuracy measurement statistics to cover this issue. One of them, and the second metric to present is the Mean Absolute Error (MAE), which is characterized by measuring the average magnitude from the absolute value of prediction errors. Its equation is:
@@
where the equation variables follow the same notations as the ones addressed for MAPE. MAE expresses the prediction performance error in absolute value; therefore, a smaller result signifies better predictive performance, while higher results represent poorer performance.
@@
We adopt as the third performance metric Root Mean Squared Error (RMSE) which captures the error value in the same value unit as the variable being predicted. Its formula is defined by
@@
where the equation variables follow the same notation as the ones explained previously. Although the definitions have value unit differences, a low error value indicates a high predictive performance in RMSE and similarly, a high error value depicts a poor predictive performance. We provide further results on AIC and BIC model selection criteria in the provided online appendix (see Section 3.7).
3.5.6 Performance evaluation
@@
The next step was to train and evaluate the performance of the selected prediction approaches using the three data sets explained in Section 3.4.2. The models were trained and tested using the Walk-Forward Train-Test validation technique [53]. This technique was chosen as it respects the temporal nature of the data and does not rely on randomness, yielding a fair comparison between methods. The Walk-Forward Train-Test technique trains and tests a model greedily and iteratively. Each iteration trains a model, predicts the value of the next data point of the time series, and compares the prediction against the real value. The real value is added consequently to the train set and the model is trained again with the newly extended train set data. This cycle is continued until all test data points are predicted. In our study, the models initially start the training with 80% of the data as training data and subsequently predict and test every new data point. Figure 2 shows a graphical representation of the approach, where i denotes the new data point in each iteration. The described model training and testing phases performed in the performance evaluation stage of this study were implemented equally in each of the adopted models, with the same testing set of data, for the sake of fairness in the predictive performance comparison.
3.6 Model Execution
@@
In this section, we emphasized the description of the model execution process of the defined TSA models to facilitate practitioners’ understanding of the results presented in Section 4 and encourage the use of the proposed models in the SE community. Nonetheless, the prediction process described in Section 3.5.6 is similarly followed for all projects with each of the considered prediction models using different data sets. Note that the ML models are tested with all three data sets, but the time-dependent models only use the biweekly and monthly time series data. After the prediction process is concluded, the results are visualized, and therefore organized, by aggregating the results provided by each model with all the collected projects for the defined three data sets (raw data, biweekly data, and monthly data). The aggregation of the results per dataset for each model combination is done by taking the average of the performances for the different projects in the respective data set as displayed in Tables 4 and 5 for instance.
3.6.1 Execution process for ARIMAX & SARIMAX models
@@
Within the context of the modelling process, both ARIMAX and SARIMAX models are similarly built in their initial stage. The last one stands as an adjusted version of the former counterpart. Therefore, for each project, we built both an ARIMAX and SARIMAX model. Given the existing lack of normality in the distribution of the independent variables (see Table 1 and Figure 1) we log-transformed their observations in order to standardize the independent variable observations. Variable transformation is a common preprocessing technique used in the field of predictive analytics due to its benefits in reducing the model complexity and augmenting the prediction performance [54].
The first step of the modelling process consisted of fitting the models with the training data and following the notions described in Section 3.5.3 for each model accordingly to perform the model parameter tuning. Likewise, we performed a variable backward selection process which consists of iteratively removing such model variables that reduce the quality of the model by assessing multiple variable combinations [29]. This process consists of the model building and evaluation of multiple combinations of the considered best independent variables. This iterative process is repeated until the variable combination that better improves the model quality is reached.
@@
Figure 3 provides a graphical representation of the seasonal decomposition process performed for the SARIMAX model with the biweekly SQALE index data of the Apache httpscore project to achieve the best model parameters. Described in a top-down structure, the first plot provides the observed progress of the SQALE index across time, the second and the third plots provide the existing trend and seasonality pattern, and the fourth one depicts the remaining residuals or noise in the data after de-trending and de-serializing has been performed. The seasonal decomposition process is conducted every time a seasonally adjusted model is built. Hence, we provide the visualization of the seasonal decomposition for each of the considered projects in the provided online appendix (see Section 3.7).
@@
For each fitted TSA model, the explained iterative process concludes after all the combinations have been explored and the best parameter order has been found. This stage provides a model that shows a good fit for the data and provides model residuals normally distributed, thus showing good model quality. We provide a graphical representation of this stage in Figure 4, where the model diagnostics are presented for the same example project as in Figure 3.
@@
The standardized residuals help to identify the existing anomalies or obvious patterns if the former ones are not centered around zero, therefore meaning that some patterns remain uncaptured by the model. The second plot provides the histogram of the model residuals where in comparison with the standard normal distribution the kernel density estimate seems to be approximately normally distributed. In the third plot, the model residuals are compared against the Q-Q plot standards, showing a deviated distribution of the residuals from normality. Finally, the correlogram or Auto Correlation Function (ACF) plot shows the residuals autocorrelations, which should not be after the model is fitted. Hence, they should not be exaggeratedly outside the marked confidence interval.
3.6.2 Execution process for ARIMA+LM and SARIMA + LM models

### v2

@@ 3.6.4 Setting the stage for Machine Learning models (RQ2 - RQ4)
In RQ2 and RQ4, we compare the prediction performance of TSA approaches with that of ML algorithms. Therefore, we also investigate the ability of ML models to predict code TD by applying a collection of linear and nonlinear ML algorithms. To run the comparison in equal conditions with all the models considered in this study, the same time series data used to train and test the TSA methods are also used to train and test the selected ML algorithms. Further descriptions of the training and performance evaluation of the models studied can be found in Section 3.6.6.
@@
The ML methods used for the comparison are selected to provide a wide predictive perspective, as each of them textualizes different aspects of the data. Likewise, the chosen ML algorithms have been extensively used in the recent literature for their ability to predict software quality characteristics, such as code TD [12, 9, 10, 54, 55].
@@
Multiple Linear Regression (MLR) is a powerful statistical model commonly used to analyze the relationship between a continuous dependent variable and a set of independent explanatory variables. Through this relationship, the model can provide insight into how changes in the independent variables affect the value of the response variable while maintaining the linearity assumption [56].
@@
Stochastic Gradient Descent (SGD) algorithm is based on an iterative decrease in prediction loss by changing the model parameters. Specifically, for each iteration, the SGD algorithm randomly selects a training set to bring randomness to the optimization problem, and the algorithm continues to optimize the result until it converges [57].
@@
Lasso regression (L1) is a linear regression technique specialized in high-dimensional explanatory variable sets. This technique performs an optimization problem where model-independent variables are penalized if they are not important to the model. Thus, through model simplification, L1 achieves linear regression with the most relevant explanatory variables [58].
@@
Ridge regression (L2) is another linear regression technique that, through a regularization term, avoids overfitting the model. The L2 technique brings a penalization for the given model parameters that denote extreme values, and thus stabilizes the model, which results in advantages for applications where the number of model-independent variables is high [59].
The selected ML models assuming nonlinearity (NL) are:
@@
Support Vector Machine (SVM) is a well-established ML algorithm used for classification and regression tasks. The regression model aims to find the most optimal hyperplane that considers the characteristics of the existing explanatory variables to explain the values of the output response variable. For this, the model performs a minimization process of the deviation between the real value and the predicted value [60]. We considered using the so-called Gaussian kernel or Radial Basis Function as the function to build the model to get the perspective of a non-linear approach.
@@
Extreme Gradient Boost (XGB) is an ML algorithm that is part of the family of gradient boost methods. Based on the ensemble learning methodology, the XGB algorithm employs multiple decision trees in a greedy format, where each of the subsequent trees corrects or refines the wrong results from the previous tree. Through this technique, the algorithm defines weights for the included independent variables that require a higher emphasis [61].
@@
Random Forest regression (RF) also belongs to the ensemble learning family and is built based on the ensemble of decision trees. Based on random sampling, the RF algorithm performs multiple decision trees and the output is the aggregated value from the results obtained in the trees performed [62].
3.6.5 Performance metrics (RQ1 - RQ5)
@@
We used three performance metrics to compare and evaluate the prediction approaches and models.
@@
The first performance metric is the Mean Absolute Percentage Error (MAPE). MAPE is a statistical metric commonly used to measure prediction precision. Quantifies the magnitude of errors between the predicted values and the actual observed values by calculating the mean of the absolute value of the prediction error. Its formula is the following.
@@
where is the number of observations, is the actual observed value, and is the i-th predicted value. As shown by the formula, MAPE is represented as a percentage metric, where a smaller percentage of error means better predictive performance, and the opposite is true with a higher percentage.
@@
MAPE has limitations when the actual observations are small or close to zero, which introduces bias into the model training. Therefore, we selected alternative accuracy measurement statistics to cover this issue. One of them, and the second metric to present is the Mean Absolute Error (MAE), which is characterized by measuring the average magnitude from the absolute value of prediction errors. Its equation is
@@
where the variables follow the same notation as the ones addressed for MAPE MAE expresses the prediction performance error in absolute value; therefore, a smaller result signifies better predictive performance, while higher results represent poorer performance.
@@
We adopt as the third performance metric root mean squared error (RMSE) which captures the error value in the same value unit as the variable being predicted. Its formula is
@@
where the variables follow the same notation as the ones explained previously. Although the definitions have value unit differences, a low error value indicates a high predictive performance in RMSE, and similarly, a high error value depicts a poor predictive performance. We provide further results on AIC and BIC model selection criteria in the online Appendix (see Section 9).
3.6.6 Performance evaluation (RQ1 - RQ5)
To evaluate the performance of the confronted models, we trained and tested the TSA and ML models on the two time-series datasets generated during the data serialization in Section 3.4.2. We adopted the Walk-Forward Train-Test validation technique [63] for the performance evaluation of all models considered in this study. This technique was chosen because it respects the temporal nature of the data and does not rely on randomness in the order of the observations. This helped to produce a fair comparison between the confronted methods, as especially ML models are not structurally designed to understand the temporal order existing in the fitted data.
@@ 3.6.6 Performance evaluation (RQ1 - RQ5)
The Walk-Forward Train-Test technique trains and tests a model greedily and iteratively. Each iteration trains a model, predicts the value of the next data point of the time series, and compares the prediction against the real value. The real value is consequently added to the train set, and the model is trained again with the newly extended train set data. This cycle continues until all test data points are predicted. In our study, the models start training with 80% of the data as training data and then predict and test every new data point. Figure 3 shows a graphical representation of the approach, where i denotes the new data point in each iteration. The described model training and testing phases performed in the performance evaluation stage of this study were implemented equally in each of the adopted models, with the same data testing set, for the sake of fairness in the predictive performance comparison.
3.6.7 Practitioners’ forecasting perceived valuable (RQ6)
This section presents the data analysis of our work. Our survey includes closed and open questions. Therefore, we select different analysis methods for the two types of survey output. To analyze the responses to the closed questions, we initially employed descriptive statistics to gain a clearer understanding of the data. For ordinal and interval data, we focused on the mode and median to assess central tendency, while for nominal data, we calculated the distribution of participants’ choices for each option.
Regarding the Likert scale question, Q12, the possible value ranged from 0 (“not at all useful”) to 10 (“extremely useful”). We present both the values as they have been reported by the interviewees and also the interpretation of the Net Promoter Score (NPS) [64]. NPS is a widely used metric for measuring customer loyalty and satisfaction. It classifies respondents into three categories based on their likelihood to recommend a product, service, or company:
Promoters (9-10): Highly satisfied and likely to advocate for the product.
Passives (7-8): Moderately satisfied but not enthusiastic enough to promote it.
Detractors (0-6): Unhappy customers who may discourage others from using the product.
The NPS is calculated as: . Therefore, a higher NPS indicates stronger customer loyalty. Hence, we employ such a score to measure the practitioner’s perceived value of our model.
Regarding open-ended questions, we employ qualitative data analysis techniques suggested by Strauss and Corbin [65] and Seaman and Yuepu [66]. Qualitative analysis assists in addressing questions such as “What is happening in this situation?” when we aim to uncover how individuals understand their experiences and manage them over time in amidst evolving conditions [41].
We adopted an inductive approach to develop a new theory based on qualitative data. The open questions for RQ6.1 and RQ6.2 were manually coded as follows. Two authors independently coded responses to related questions adopting thematic analysis [65, 27]. We addressed disagreement through discussion, and codes were organized into a hierarchy of benefits and limitations until saturation was reached.
3.7 Model Execution (RQ1 - RQ5)
@@
In this section, we textualize the description of the model execution process of the defined TSA models to facilitate practitioners’ understanding of the results presented in Section 4 and encourage the use of the proposed models in the SE community. However, the prediction process described in Section 3.6.6 is followed similarly for all projects with each of the prediction models considered using different data sets. After the prediction process is completed, the results are visualized and, therefore, organized by aggregating the results provided by each model with all the collected projects for the two defined data sets (biweekly data and monthly data). The aggregation of the results per data set for each model combination is done by taking the average performance of the different projects in the respective data set as displayed in Table 5 and Table 6 for instance.
3.7.1 ARIMAX & SARIMAX models (RQ1 - RQ4)
@@
Within the context of the modeling process, both the ARIMAX and SARIMAX models are built similarly in their initial stages. The last one stands as an adjusted version of the former counterpart. Therefore, for each project, we built both an ARIMAX model and a SARIMAX model. We computed the exploratory analysis of the model variables to examine their distributional characteristics (see Table 1 and Figure 1). All model variables denoted high skewness values, which was visually represented in the displayed distribution boxplots. Therefore, we applied the log-transform technique on the independent variables of the model to standardize their distribution and thus reduce the complexity of the model variables during the learning process. Variable transformation is a common preprocessing technique used in the field of predictive analytics due to its benefits in reducing the complexity of the model and increasing the prediction performance [67].
The first step of the modeling process consisted of leveraging the Backward variable selection technique [33]. The stepwise backward procedure begins with fitting the model with the set of important independent variables initially considered. Subsequently, it sequentially assesses the model and removes variables from the model to train it back. In each sequence, it selects the variable whose removal improves the goodness of fit of the model. The process stops when any further removal leads to a poorer model fit, therefore selecting the set of independent variables that provides the most optimal model results. To assess the impact of each combination of variables accordingly and following the concepts described in Section 3.6.3, we performed the model parameter tuning approach with the Auto-arima algorithm to obtain the results of the criteria AIC and BIC. We provide a concise description of the parameter tuning process in appendix A.
@@
Figure 4 provides a graphical representation of the seasonal decomposition process performed for the SARIMAX model with the biweekly SQALE index data of the Apache httpcore project to achieve the best model parameters Described in a top-down structure, the first plot provides the observed progress of the SQALE index across time, the second and the third plots provide the existing trend and seasonality pattern, and the fourth one depicts the remaining residuals or noise in the data after de-trending and de-serializing has been performed The seasonal decomposition process is conducted every time a seasonally adjusted model is built Hence, we provide the visualization of the seasonal decomposition for each of the considered projects in the provided online appendix (see Section 9).
@@
For each of the studied software projects, we performed the multivariate TSA model fitting following the explained stepwise process. After each process, the model would provide the best goodness-of-fit results for the fitted data, and normally distributed model residuals, therefore showing good model quality. We provide a graphical representation of this stage in Figure 5, where the model diagnostics are presented for the same example project as in Figure 4.
@@
The standardized residuals help to identify the existing anomalies or obvious patterns if the former ones are not centered around zero, therefore meaning that some patterns remain uncaptured by the model. The second plot provides the histogram of the model residuals, where in comparison with the standard normal distribution the kernel density estimate seems to be approximately normally distributed. In the third plot, the model residuals are compared against the Q-Q plot standards, showing a deviated distribution of the residuals from normality. The correlogram or Auto Correlation Function (ACF) plot shows the residuals autocorrelations, which should not be present after the model is fitted. Hence, they should not be exaggerated outside the marked confidence interval.
3.7.2 ARIMA+LM and SARIMA + LM models (RQ1 - RQ4)